OpenStudy (anonymous):

what is the solution of the proportion? 14/12=d/48

OpenStudy (anonymous):

do you know about cross-multiplying? when you have a proportion like (a/b) = (c/d) you can cross multiply to get (a)(d) = (b)(c)

OpenStudy (anonymous):

yes which would it be 672 or 56?

OpenStudy (anonymous):

The answer is 56. You multiply 48 by 14 and divide by 12.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

48 * 14 = 672 Then you need to divide 672 by 12

OpenStudy (anonymous):

which gives you 56? thank you both